For the diprotic weak acid H2A, ?a1=2.9×10−6 and ?a2=8.4×10−9

What is the pH of a 0.07000.0700 M solution of H2A?

What are the equilibrium concentrations of H2A and A2− in this solution?
Answer #1

Answer:-

1. DH = 3.35.

2. equilibrium [H2A] = 0.0695409 M

3. equilibrium [A2-] = 8.4 x 10-9M

Explanation :-

Given:-

MolarityOf HA=0.0700M

Kal = 2.9 x 10-6

K2 = 8.4 x 10-9

Now, First dissociation is

HA=H+ + HA ...(1)

The ICE Table is

H_{2}A \rightleftharpoons H^{+} + HA^{-}
initial 0.07 0 0
change -x +x +x
equilibrium 0.07 -x x x

Now, from equilibrium (1) we have,

[H+][HA] Kal = ! (H2A

2.9 x 10-6 = 0,07 -

but, x << 0.07 hence, 007-7007

therefore,

2.9 \times 10^{-6} = \frac{x^{2}}{0.07}

ie. x^{2} = 2.9 \times 10^{-6} \times 0.07

x^{2} = 0.203 \times 10^{-6}

x = 0.451 \times 10^{-3} = 4.51 \times 10^{-4}

ie. [H+] = [HA-] = 4.51 \times 10^{-4}

Now second dissociation is

HA^{-} \rightleftharpoons H^{+} + A^{2-}...(2)

The ICE table is

HA^{-}   \rightleftharpoons H^{+} + A^{2-}
initial 4.51 \times 10^{-4} 4.51 \times 10^{-4} 0
change -x 4.51 \times 10^{-4} + x +x
equilibrium 4.51 \times 10^{-4} - x 4.51 \times 10^{-4} + x x

Now from equilibrium (2) we have,

K_{a2} = \frac{[H^{+}][A^{2-}]}{[HA^{-}]}

ie. 8.4 \times 10^{-9} = \frac{(4.51 \times 10^{-4} +x)(x)}{(4.51 \times 10^{-4} - x)}

but x is very small

therefore,

8.4 \times 10^{-9} = \frac{(4.51 \times 10^{-4})(x)}{(4.51 \times 10^{-4})}  

ie. 8.4 \times 10^{-9} = x

Therefore, [H+] = (4.51 \times 10^{-4} + 8.4 \times 10^{-9}) \approx 4.51 \times 10^{-4} M

Now,

p^{H} = - log [H^{+}]

pH = -log(4.51 x 10-4)

p^{H} = - (-3.35)

DH = 3.35

Again,

equilibrium [H_{2}A] = 0.07 - x

equilibrium [H_{2}A] = 0.07 - 4.51 \times 10^{-4} = 0.069549 M  

and

equilibrium [A2-] = 8.4 x 10-9M